OpenStudy (anonymous):

If X is b(n,p) show that E[((X/n)-p)^2]= (p(1-p))/n

OpenStudy (anonymous):

show that \[E[(\frac{ X }{ n } - p)^2 =\frac{ p(1-p) }{ n }\]